TMS Meeting Minutes, September 9, 2021

Attendees: Andrew, Breann, Heejoo, John, Olga, Trenton, Yang (Zoom Meeting)

Recent Updates on TMS TCP/IP:
  • 9/3 TCP/IP data acquisition code works well
  • Threshold filtering is too aggressive in filtering results
  • On this particular night, we were using TMS passively while collecting superfoc sequences
  • The large focus adjustments caused measurements to exceed the std. dev. threshold, so many measurements were dropped
  • John updated the TMS code to prevent the system from collecting data while the telescope is slewing
  • John notes that there were instances of glitches that actually snuck past the std. dev. filter when they most certainly should not have
  • Heejoo to take a closer look to see if this was associated with a specific channel number
Laser Spectrum Error:
  • We encountered an error which we do not yet understand the cause of yet
  • Apparently, the laser 2 spectrum was cut off halfway, and Mark had never seen this on any other system
  • The system was only finding half the number of peaks needed on the laser spectrum, as the spectrum was shifted halfway off the plot
  • To solve this problem, power cycling the MultiLine machine seems to fix the issue
  • Power cycling only the PC is not sufficient; the whole MultiLine unit must be power cycled
  • For the time being, it will be a member of the TMS team who will monitor for this spectrum error
  • At the time of the error, John noted that the TCP/IP control interface only shows a timeout on the measurement with no other error message
  • As it stands, all kinds of errors can be combated by power cycling the MultiLine system
TMS Opportunity Tomorrow Night:
  • Heejoo would like to check again with TMS running in passive mode
  • There is possibly an opportunity tomorrow night after the ARGOS testing is finished up
  • John to ask for time from 3:45 - 4:30 PM MST
TMS Fiber Problems:
  • One of the fiber connectors at the MultiLine system is broken
  • Heejoo and Trenton found that the fiber connector may be reassembled with the correct rotation to achieve a 20% signal
  • All fiber bundles are hanging in the TMS cabinet to minimize the load on the individual fibers
  • Heejoo asked Leroy to find the fiber inspection tool so that we may test the fibers for signal fidelity
  • Heejoo is concerned about the effects of routing fibers on moving parts
  • Currently, the Sx fiber bundle routes using a plastic PVC pipe
  • The Dx side is routed as a bare bundle of fiber without any covering
  • As of right now, we have a lot of active channels on the Sx side and fewer on the Dx side
TMS Collimator Cap Development:
  • We would like to have a cap to attach to the front of the collimator to improve collimator performance in the visible
  • Andrew suggests having a negative lens to introduce beam divergence, but to focus on emphasizing transmission in the 1532 nm IR band
  • The signal in the IR will be reduced by a significant fraction while the visible light will be vastly improved
Thermal Aberration Analysis:
  • Andrew sent out a new document on thermal analysis, where section 3 is the new material
  • Andrew notes some "bad" channels from the faceplate thermocouples, but these data values are still usable, even with the added 10000 K
  • Analysis focus is on the relative temperatures between faceplate and backplate
  • 21000 x Delta T gives the temperature-induced Z4
  • There appears to be a telemetry issue, where thermocouple data is being reported with the same value two or three times in a row
  • John to look into this and see why the thermocouple data is not updating at each timestep
  • For seeing values less than or equal to 0.8, the convergence criterion is 250 nm for FPIA
  • The model appears to be doing work, but in some cases it is not enough to keep us collimated
  • The biggest errors appear to occur with the largest thermal changes
  • Trenton developing a 3D FEA model to help refine the thermal model further, going further than what RFCML has done
  • Olga wonders about the errors in the crosstalk terms between Z4, Z11, and Z22 in the collimation model and their effect on this thermal model
Engineering Time:
  • There are three engineering nights left, Friday, Saturday, and Sunday
  • We would like to take advantage of these times, if possible
  • Olga notes that the checkout schedule might actually end on 9/15, not 9/12
  • If this is the case, we should have lots of time to test TMS

-- TrentonBrendel - 09 Sep 2021
Topic revision: r1 - 09 Sep 2021, TrentonBrendel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback